OpenStudy (anonymous):

The area of a football field is x2 + 3x – 10. Find the length and width of the field in terms of x.

OpenStudy (ash2326):

The area of a football field is \[x^2+3x-10\] Let's find factors of this, one factor is length other is width \[x^2+3x-10\] Note that the coefficient of x^2 is 1 and the term without x is -10 We have to find factors of \(1\times -10=-10\) such that their sum is +3 5 and -2 satisfy this \[(5)+(-2)=3\] and \[5 \times -2=10\] so 3x can be written as \[3x=5x-2x\] we have \[x^2+5x-2x-10\] take x common from the first two terms and -2 from the last two terms \[x(x+5)-2(x+5)\] now take (x+5) common from the terms \[(x+5)(x-2)\] So we have found the two factors (x+5) is the length (x-2) is the width \[x(

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

The area of a football field is x2+3x−10 Let's find factors of this, one factor is length other is width x2+3x−10 Note that the coefficient of x^2 is 1 and the term without x is -10 We have to find factors of 1×−10=−10 such that their sum is +3 5 and -2 satisfy this (5)+(−2)=3 and 5×−2=10 so 3x can be written as 3x=5x−2x we have x2+5x−2x−10 take x common from the first two terms and -2 from the last two terms x(x+5)−2(x+5) now take (x+5) common from the terms (x+5)(x−2) So we have found the two factors (x+5) is the length (x-2) is the width \[x(